OpenStudy (anonymous):

The National Association for Women in Science asked recent high school grads if they had taken certain science classes. Of those surveyed, 90 said they had taken physics, 138 said they had taken chemistry and 45 said they had taken both. Thirty-two said they had taken neither. How many recent high school grads were surveyed?

OpenStudy (anonymous):

what i did was add the numbers up and got an answer. would that be correct or no?

OpenStudy (anonymous):

If you're familiar with set notation:\[P \cup C=C+P-CnP=90+138-45=183\]I used n as intersection.

OpenStudy (shadowfiend):

Hm. Just realized this isn't in our equation editor -> you can use `\cap' for intersection (where \cup is union).

OpenStudy (anonymous):

Great! Thanks.